OpenStudy (anonymous):

A farmer plans to create a rectangular garden that he will enclose with chicken wire. The garden can be no more than 30 ft wide. The farmer would like to use at most 180 ft. of chicken wire. Write a system of linear inequalities that models this situation.

OpenStudy (anonymous):

@jim_thompson5910, Can you help me

jimthompson5910 (jim_thompson5910):

Let L = length and W = width

jimthompson5910 (jim_thompson5910):

"The garden can be no more than 30 ft wide" what does that translate into?

OpenStudy (anonymous):

30 + x = w

OpenStudy (anonymous):

Is that right @jim_thompson5910

OpenStudy (anonymous):

30 + x =w @jim_thompson5910

jimthompson5910 (jim_thompson5910):

"The garden can be no more than 30 ft wide" means that \(\Large w \le 30\)

OpenStudy (anonymous):

@jim_thompson5910. Is that the answer

jimthompson5910 (jim_thompson5910):

no, that's one part of the answer

jimthompson5910 (jim_thompson5910):

"Write a system of linear inequalities" implies there is more than one inequality

OpenStudy (anonymous):

Oh so how would i find the other part

jimthompson5910 (jim_thompson5910):

how would translate "The farmer would like to use at most 180 ft. of chicken wire" ?

OpenStudy (anonymous):

F less than 180 ?

jimthompson5910 (jim_thompson5910):

Say P is the perimeter, so if "The farmer would like to use at most 180 ft. of chicken wire", then \(\Large P \le 180\) The perimeter P is P = 2L + 2W, so we can replace the P in the inequality with 2L + 2W to get \(\Large 2L + 2W \le 180\) \(\Large 2(L + W) \le 180\) \(\Large L + W \le \frac{180}{2}\) \(\Large L + W \le 90\) see how I'm getting all this?

OpenStudy (anonymous):

Yes...

jimthompson5910 (jim_thompson5910):

any questions on it?

jimthompson5910 (jim_thompson5910):

i noticed you were typing, but you stopped

OpenStudy (anonymous):

Sorry my computer died. Now im using my phone. Yes i get it. And would those other ones work for the question @jim

jimthompson5910 (jim_thompson5910):

The final answer would be \(\Large W \le 30\) \(\Large L + W \le 90\) \(\Large L > 0\) \(\Large W > 0\) I'm adding those last two inequalities because I'm forcing the length and width to be positive numbers

jimthompson5910 (jim_thompson5910):

So the final answer is a combination of what is discussed above

OpenStudy (anonymous):

@jim_thompson5910, Thank you so much, i can not thank you enough!

jimthompson5910 (jim_thompson5910):

you're welcome